# Demonstration of a parity–time symmetry breaking phase transition using superconducting and trapped-ion qutrits

### v2

@@ I Introduction
Quantum simulation is one of the key prospective applications for quantum computing Lloyd (1996); Buluta and Nori (2009); Cirac and Zoller (2012); Georgescu et al. (2014); Fedorov et al. (2022); Hoefler et al. (2023). It uses a well-controlled quantum device to replicate the behavior of the system of interest. There are two main approaches to quantum simulation. One is the analog quantum simulation, which relies on special-purpose quantum systems and can be based on a variety of platforms including superconducting transmons Houck et al. (2012); Hartmann (2016), trapped ions Monroe et al. (2021); Blatt and Roos (2012), neural atoms Browaeys and Lahaye (2020); Gross and Bloch (2017), and photons Aspuru-Guzik and Walther (2012); Hartmann (2016). These systems have been used to study non-trivial quantum effects Daley et al. (2022); Bloch (2005); Bloch et al. (2008, 2012); Blatt and Roos (2012); Kjaergaard et al. (2020), e.g. reproducing phase transitions in quantum many-body systems Bloch (2005); Bloch et al. (2008); Blatt and Roos (2012); Bernien et al. (2017); Keesling et al. (2019). While the analog simulators have arguably reached the practical quantum advantage threshold, the scope of their applications is likely to remain limited to a class of models that can be simulated and the level of precision in quantitative predictions Daley et al. (2022). Another approach is to use digital quantum devices Lloyd (1996) capable of universal quantum computation and in principle not limited in the type of systems they can describe Martinez et al. (2016); Lanyon et al. (2011). Digital quantum simulation can address various physical Bassman et al. (2021); Bauer et al. (2020); Georgescu et al. (2014); Barends et al. (2015) and chemical Cao et al. (2019); McArdle et al. (2020a, b) problems intractable for classical computing. However, reaching sufficient precision in quantitative predictions calls for significant improvements in the quantum hardware, and likely requires fault-tolerance von Burg et al. (2021). Modern quantum computing devices are designed to perform reversible operations and natively support only unitary gates Bennett and DiVincenzo (2000). Simulation of standard Hermitian Hamiltonians fits well within this framework Buluta and Nori (2009); Cirac and Zoller (2012); Georgescu et al. (2014), yet modeling the behavior of non-conservative quantum systems is equally valuable. Understanding Markovian and non-Markovian dynamics of open quantum systems Rivas and Huelga (2011); Lidar (2019); Ashida et al. (2020); Luchnikov et al. (2022a) is important to describe a range of physical phenomena, such as decoherence Schlosshauer (2019), thermalization D’Alessio et al. (2015); Nandkishore and Huse (2015); Reichental et al. (2017); Žnidarič et al. (2010); Abanin et al. (2019), noise characterization Harper et al. (2020); Youssry et al. (2020); Georgopoulos et al. (2021), and others as well as for realizing quantum control protocols  James (2021); Dong and Petersen (2009); Luchnikov et al. (2022b). It is in fact possible to simulate non-unitary dynamics using a reversible quantum computer, and numerous techniques have been developed to this end, including methods based on linear combination of unitaries Wei et al. (2016); Schlimgen et al. (2021); Zheng (2021) or dilation Hu et al. (2019); Head-Marsden et al. (2021); Hu et al. (2022); Sweke et al. (2016). Effectively, the Hilbert space can be split into two parts – one part encoding the system of interest, and the other the environment. An interaction between the system and the environment is then simulated by a properly engineered unitary evolution of the total system.
@@ II $\mathcal{P}\mathcal{T}$-symmetric systems and simulation
A $\mathcal{P}\mathcal{T}$-symmetric system is governed by a non-Hermitian Hamiltonian $H$, which is invariant with respect to the combined parity $\mathcal{P}$ and time-reversal $\mathcal{T}$ transformations, $[H,\mathcal{P}\mathcal{T}]=0$. The characteristic polynomial of a $\mathcal{P}\mathcal{T}$-symmetric Hamiltonian is always real, and hence the eigenvalues are either all real or come in complex-conjugate pairs. In the former case, the system is said to be in the $\mathcal{P}\mathcal{T}$-unbroken phase. The regime with complex eigenvalues corresponds to the $\mathcal{P}\mathcal{T}$-broken phase and typically arises as the gain and loss terms become sufficiently strong, so that the non-unitary aspects of the dynamics dominate the internal interactions Bender et al. (2018). $\mathcal{P}\mathcal{T}$-symmetric systems feature many unusual properties such as complex spectrum, exceptional points and coalescence of eigenvectors, non-conservation of the trance distance between quantum states, and breaking entanglement monotonicity. In this work, we focus on probing the $\mathcal{P}\mathcal{T}$-symmetry breaking phase transition, and the associated qualitative change in the dynamics.
@@ II.1 Two-level $\mathcal{P}\mathcal{T}$-symmetric system
We henceforth refer to $H$ simply as the $\mathcal{P}\mathcal{T}$-symmetric Hamiltonian. The real parameter $r$ quantifies the strength of the gain and loss (diagonal) terms compared to the inter-level interactions. The parity operator is $\mathcal{P}=\sigma_{x}$, and the time-reversal operator acts by complex conjugation $\mathcal{T}(\cdot)=(\cdot)^{*}$. The Hamiltonian (3) is $\mathcal{P}\mathcal{T}$ symmetric for any real value of $r$, i.e. $[H,\mathcal{P}\mathcal{T}]=0$. The eigenvalues of $H$ are $h_{\pm}=\pm h,h=\sqrt{1-r^{2}}\ .$ For $r<1$ the eigenvalues are real and $\mathcal{P}\mathcal{T}$-symmetry is unbroken, while $r>1$ leads to purely imaginary values of $h$ and hence breaks $\mathcal{P}\mathcal{T}$-symmetry. The value $r=1$ corresponds to the exceptional point. Similarly to the unitary case, in the $\mathcal{P}\mathcal{T}$-unbroken phase eigenvectors of the system $|\psi_{\pm}\rangle$ acquire complex phases during the time evolution, and level populations manifest Rabi-like oscillations. As the phase transition point $r=1$ is approached from within the $\mathcal{P}\mathcal{T}$-unbroken phase, the period of oscillations $T=\frac{2\pi}{\sqrt{1-r^{2}}}$ grows and diverges at $r=1$. After that, the $\mathcal{P}\mathcal{T}$-broken regime with $r>1$ exhibits an exponential relaxation to the ground state with time $\tau=\frac{1}{\sqrt{r^{2}-1}}$, without the oscillatory behavior Kawabata et al. (2017b). Our main goal in this work is to probe this expected transition experimentally.
@@ II.2 Embedding non-Hermitian evolution into a unitary operator
Here $\lambda$ is a scalar factor and $B,C,D$ are arbitrary matrix entries compatible with the unitarity of $U(t)$. Such embeddings arise in many settings. Stinespring dilation of CPTP maps is an example Stinespring (1955). In the context of quantum algorithms based on transformations of singular values, they are known as block encodings Gilyén et al. (2018); Martyn et al. (2021). An arbitrary operator $V(t)$ can be represented in the form of Eq. (6), as long as its operator norm satisfies $||V(t)||\leq 1$ (for details see Appendix A). Operators with larger norms can be embedded if rescaled appropriately $V(t)\to V(t)/\lambda$, as we indicated in Eq. (6). To apply the evolution operator $V(t)$ to an arbitrary initial state $|\psi\rangle$, one embeds $|\psi\rangle$ into the larger space and applies $U(t)$ to the result
@@
To probe the structure of the embedded state $V(t)|\psi\rangle$, the measurements are performed on the full resulting state, and the outcomes lying in the correct subspace are post-selected. The success probability of the post-selection is equal to $\lambda^{-2}\langle\psi|V(t)^{\dagger}V(t)|\psi\rangle$. Hence, it decreases as $\lambda$ grows. From this point of view, it is optimal to choose the minimal $\lambda$ compatible with the restriction $\lambda^{-1}||V(t)||\leq 1$, which is solved by $\lambda(t)=\sigma_{\rm max}(t)$, with $\sigma_{\rm max}(t)$ being the largest singular value of $V(t)$. We note that rescaling the evolution operator by the scalar factor $\lambda(t)$ is equivalent to shifting the Hamiltonian by a time-dependent constant
@@ II.3 Simulating two-level $\mathcal{P}\mathcal{T}$-symmetric system with a unitary qutrit
The previous section contains a general discussion of embedding a non-unitary evolution operator into larger unitary dynamics. Here we consider the case where the evolution operator is that of the $\mathcal{P}\mathcal{T}$-symmetric qubit Eq. (3), while the embedding system is a qutrit. There is an additional subtlety in this case, stemming from the fact that a single auxiliary dimension is not sufficient to simulate a general operator $V(t)$. However, precisely for the case when the scalar factor is chosen to be $\lambda(t)=\sigma_{\rm max}(t)$ the embedding is possible, see Appendix A. The evolution operator $V(t)$ for the $\mathcal{P}\mathcal{T}$-symmetric qubit (3) can be written as $V(t)=\cos(ht)-i\frac{\sin(ht)}{h}H,$ and its singular values read
@@ III Demonstration with trapped-ion qutrits
Here we report the simulation using a trapped ion quantum processor, which is an upgraded version of the recently presented setup (see Refs. Aksenov et al. (2023); Zalivako et al. (2023)). It is a chain of ten ${}^{171}$${}^{+}$ ions inside a linear Paul trap. Qudits are encoded in Zeeman sublevels of ${}^{2}S_{1/2}(F=0)$ and ${}^{2}D_{3/2}(F=2)$, with the qudit dimension up to $d=6$. In this work we employ only three states of each qudit, which we further refer as $|0\rangle=\,^{2}S_{1/2}(F=0,m_{F}=0)$, $|1\rangle=\,^{2}D_{3/2}(F=2,m_{F}=0)$ and $|2\rangle=\,^{2}D_{3/2}(F=2,m_{F}=1)$. More details on the experimental setup are given in the  Appendix B. Information about initialization, quantum gates, and readout procedures are also given there.
@@ Acknowledgements
We acknowledge Alexey Ustinov for fruitful discussions and comments on the manuscript. Authors acknowledge support by the Federal Academic Leadership Program Priority 2030, National University of Science and Technology MISIS: Strategic Project Quantum Internet - A.S.K., I.A.S., A.V.K., E.Yu.E., N.N.A., and N.M.; the work of N.A.N, A.S.N., and A.K.F. has been supported by Project No. K1-2022-027 at MISIS. N.A.N. thanks the support of the Russian Science Foundation Grant No. 23-71-01095 (theoretical modeling of quantum circuits). E.O.K. thanks for the support of the Russian Science Foundation Grant No. 19-71-10091 (study of embedding qubits’ non-Hermitian dynamics into qutrits). I.V.Z, A.S.B., K.P.G., N.V.S., A.E.K., K.Yu.K., I.A.S., and N.N.K acknowledge the support of the Russian Roadmap on Quantum Computing (Contract No. 868-1.3-15/15-2021, October 5, 2021) in the development of the trapped-ion processor. The research was carried out within the framework of Russian Roadmap on Quantum Computing (Contract No. 151/21-503, December 21, 2021 dated 12/21/2021). The transmon-based device was fabricated using the equipment of MIPT Shared Facilities Center.
@@ A.1 General
fulfills the last equation in Eq. (34). Thus, the key question is whether the first equation in Eq. (34) has a solution. In fact it does, provided
@@ A.3 Gate-level implementation
The last form emphasizes that the top left block of $R_{X}^{(12)}(\theta)$ reproduces $\Sigma/\sigma_{\rm max}$, while the unspecified entries *** do not affect the resulting block encoding. They can be chosen arbitrarily (subject to the unitarity constraint), and $R_{X}^{(12)}(\theta)$ provides perhaps the simplest such choice.
@@ B.1 Initialization
Figure 6: Level scheme of the ${}^{171}$${}^{+}$ ion. Solid lines show laser fields. Dashed lines show laser fields obtained by lasers phase modulation.
@@ B.3 Readout
The readout error sources include optical pumping from the $|0\rangle$ state to the ${}^{2}D_{3/2}(F=2)$ manifold during the transient stage in the beginning of each readout cycle, non-resonant pumping between qudit states, single-qudit gates infidelities and spontaneous decay of the ${}^{2}D_{3/2}(F=2)$ states Semenin et al. (2021). All these error contributions can be rather straightforwardly reduced by further optimization of the system parameters. For instance, it was demonstrated, that a duration of a single readout cycle in multiparticle processors based on ${}^{171}$${}^{+}$ ions can be significantly decreased in comparison to our current results by reducing the amount of the stray light on the detector and increasing a photon collection efficiency Debnath (2016). This would significantly reduce errors due to spontaneous decay and non-resonant pumping.
@@ C.2.3 Readout
where $\langle$.$\rangle$ . ⟩ stands for the averaging over all trajectories and $*{}^{*}$ denotes the complex conjugation operation.

### v3

@@ I Introduction
Quantum simulation is one of the key prospective applications for quantum computing Lloyd (1996); Buluta and Nori (2009); Cirac and Zoller (2012); Georgescu et al. (2014); Fedorov et al. (2022); Hoefler et al. (2023). It uses a well-controlled quantum device to replicate the behavior of the system of interest. There are two main approaches to quantum simulation. One is the analog quantum simulation, which relies on special-purpose quantum systems and can be based on a variety of platforms including superconducting transmons Houck et al. (2012); Hartmann (2016), trapped ions Monroe et al. (2021); Blatt and Roos (2012), neural atoms Browaeys and Lahaye (2020); Gross and Bloch (2017), and photons Aspuru-Guzik and Walther (2012); Hartmann (2016). These systems have been used to study non-trivial quantum effects Daley et al. (2022); Bloch (2005); Bloch et al. (2008, 2012); Blatt and Roos (2012); Kjaergaard et al. (2020), e.g. reproducing phase transitions in quantum many-body systems Bloch (2005); Bloch et al. (2008); Blatt and Roos (2012); Bernien et al. (2017); Keesling et al. (2019). While the analog simulators have arguably reached the practical quantum advantage threshold, the scope of their applications is likely to remain limited to a class of models that can be simulated and the level of precision in quantitative predictions Daley et al. (2022). Another approach is to use digital quantum devices Lloyd (1996) capable of universal quantum computation and in principle not limited in the type of systems they can describe Martinez et al. (2016); Lanyon et al. (2011). Digital quantum simulation can address various physical Bassman et al. (2021); Bauer et al. (2020); Georgescu et al. (2014); Barends et al. (2015) and chemical Cao et al. (2019); McArdle et al. (2020a, b) problems intractable for classical computing. However, reaching sufficient precision in quantitative predictions calls for significant improvements in the quantum hardware, and likely requires fault-tolerance von Burg et al. (2021).
Modern quantum computing devices are designed to perform reversible operations and natively support only unitary gates Bennett and DiVincenzo (2000). Simulation of standard Hermitian Hamiltonians fits well within this framework Buluta and Nori (2009); Cirac and Zoller (2012); Georgescu et al. (2014), yet modeling the behavior of non-conservative quantum systems is equally valuable. Understanding Markovian and non-Markovian dynamics of open quantum systems Rivas and Huelga (2011); Lidar (2019); Ashida et al. (2020); Luchnikov et al. (2022a) is important to describe a range of physical phenomena, such as decoherence Schlosshauer (2019), thermalization D’Alessio et al. (2015); Nandkishore and Huse (2015); Reichental et al. (2017); Žnidarič et al. (2010); Abanin et al. (2019), noise characterization Harper et al. (2020); Youssry et al. (2020); Georgopoulos et al. (2021), and others as well as for realizing quantum control protocols  James (2021); Dong and Petersen (2009); Luchnikov et al. (2022b). It is in fact possible to simulate non-unitary dynamics using a reversible quantum computer, and numerous techniques have been developed to this end, including methods based on linear combination of unitaries Wei et al. (2016); Schlimgen et al. (2021); Zheng (2021) or dilation Hu et al. (2019); Head-Marsden et al. (2021); Hu et al. (2022); Sweke et al. (2016). Effectively, the Hilbert space can be split into two parts – one part encoding the system of interest, and the other the environment. An interaction between the system and the environment is then simulated by a properly engineered unitary evolution of the total system.
@@ II $\mathcal{P}\mathcal{T}$-symmetric systems and simulation
A $\mathcal{P}\mathcal{T}$-symmetric system is governed by a non-Hermitian Hamiltonian $H$, which is invariant with respect to the combined parity $\mathcal{P}$ and time-reversal $\mathcal{T}$ transformations, $[H,\mathcal{P}\mathcal{T}]=0$. The characteristic polynomial of a $\mathcal{P}\mathcal{T}$-symmetric Hamiltonian is always real, and hence the eigenvalues are either all real or come in complex-conjugate pairs. In the former case, the system is said to be in the $\mathcal{P}\mathcal{T}$-unbroken phase. The regime with complex eigenvalues corresponds to the $\mathcal{P}\mathcal{T}$-broken phase and typically arises as the gain and loss terms become sufficiently strong, so that the non-unitary aspects of the dynamics dominate the internal interactions Bender et al. (2018).
$\mathcal{P}\mathcal{T}$-symmetric systems feature many unusual properties such as complex spectrum, exceptional points and coalescence of eigenvectors, non-conservation of the trance distance between quantum states, and breaking entanglement monotonicity. In this work, we focus on probing the $\mathcal{P}\mathcal{T}$-symmetry breaking phase transition, and the associated qualitative change in the dynamics.
@@ II.1 Two-level $\mathcal{P}\mathcal{T}$-symmetric system
We henceforth refer to $H$ simply as the $\mathcal{P}\mathcal{T}$-symmetric Hamiltonian. The real parameter $r$ quantifies the strength of the gain and loss (diagonal) terms compared to the inter-level interactions. The parity operator is $\mathcal{P}=\sigma_{x}$, and the time-reversal operator acts by complex conjugation $\mathcal{T}(\cdot)=(\cdot)^{*}$. The Hamiltonian (1) is $\mathcal{P}\mathcal{T}$ symmetric for any real value of $r$, i.e. $[H,\mathcal{P}\mathcal{T}]=0$.
The eigenvalues of $H$ are $h_{\pm}=\pm h,h=\sqrt{1-r^{2}}\ .$ For $r<1$ the eigenvalues are real and $\mathcal{P}\mathcal{T}$-symmetry is unbroken, while $r>1$ leads to purely imaginary values of $h$ and hence breaks $\mathcal{P}\mathcal{T}$-symmetry. The value $r=1$ corresponds to the exceptional point.
Similarly to the unitary case, in the $\mathcal{P}\mathcal{T}$-unbroken phase eigenvectors of the system $|\psi_{\pm}\rangle$ acquire complex phases during the time evolution, and level populations manifest Rabi-like oscillations. As the phase transition point $r=1$ is approached from within the $\mathcal{P}\mathcal{T}$-unbroken phase, the period of oscillations $T=\frac{2\pi}{\sqrt{1-r^{2}}}$ grows and diverges at $r=1$. After that, the $\mathcal{P}\mathcal{T}$-broken regime with $r>1$ exhibits an exponential relaxation to the ground state with time $\tau=\frac{1}{\sqrt{r^{2}-1}}$, without the oscillatory behavior Kawabata et al. (2017b). Our main goal in this work is to probe this expected transition experimentally.
@@ II.2 Embedding non-Hermitian evolution into a unitary operator
Here $\lambda$ is a scalar factor and $B,C,D$ are arbitrary matrix entries compatible with the unitarity of $U(t)$. Such embeddings arise in many settings. Stinespring dilation of CPTP maps is an example Stinespring (1955). In the context of quantum algorithms based on transformations of singular values, they are known as block encodings Gilyén et al. (2018); Martyn et al. (2021). An arbitrary operator $V(t)$ can be represented in the form of Eq. (2), as long as its operator norm satisfies $||V(t)||\leq 1$ (for details see Appendix A). Operators with larger norms can be embedded if rescaled appropriately $V(t)\to V(t)/\lambda$, as we indicated in Eq. (2).
To apply the evolution operator $V(t)$ to an arbitrary initial state $|\psi\rangle$, one embeds $|\psi\rangle$ into the larger space and applies $U(t)$ to the result
@@
To probe the structure of the embedded state $V(t)|\psi\rangle$, the measurements are performed on the full resulting state, and the outcomes lying in the correct subspace are post-selected.
The success probability of the post-selection is equal to $\lambda^{-2}\langle\psi|V(t)^{\dagger}V(t)|\psi\rangle$. Hence, it decreases as $\lambda$ grows. From this point of view, it is optimal to choose the minimal $\lambda$ compatible with the restriction $\lambda^{-1}||V(t)||\leq 1$, which is solved by $\lambda(t)=\sigma_{\rm max}(t)$, with $\sigma_{\rm max}(t)$ being the largest singular value of $V(t)$. We note that rescaling the evolution operator by the scalar factor $\lambda(t)$ is equivalent to shifting the Hamiltonian by a time-dependent constant
@@ II.3 Simulating two-level $\mathcal{P}\mathcal{T}$-symmetric system with a unitary qutrit
The previous section contains a general discussion of embedding a non-unitary evolution operator into larger unitary dynamics. Here we consider the case where the evolution operator is that of the $\mathcal{P}\mathcal{T}$-symmetric qubit Eq. (1), while the embedding system is a qutrit. There is an additional subtlety in this case, stemming from the fact that a single auxiliary dimension is not sufficient to simulate a general operator $V(t)$. However, precisely for the case when the scalar factor is chosen to be $\lambda(t)=\sigma_{\rm max}(t)$ the embedding is possible, see Appendix A.
The evolution operator $V(t)$ for the $\mathcal{P}\mathcal{T}$-symmetric qubit (1) can be written as $V(t)=\cos(ht)-i\frac{\sin(ht)}{h}H,$ and its singular values read
@@ III Demonstration with trapped-ion qutrits
Here we report the simulation using a trapped ion quantum processor, which is an upgraded version of the recently presented setup (see Refs. Aksenov et al. (2023); Zalivako et al. (2023)). It is a chain of ten 171Yb+ ions inside a linear Paul trap. Qudits are encoded in Zeeman sublevels of ${}^{2}S_{1/2}(F=0)$ and ${}^{2}D_{3/2}(F=2)$, with the qudit dimension up to $d=6$. In this work we employ only three states of each qudit, which we further refer as $|0\rangle=\,^{2}S_{1/2}(F=0,m_{F}=0)$, $|1\rangle=\,^{2}D_{3/2}(F=2,m_{F}=0)$ and $|2\rangle=\,^{2}D_{3/2}(F=2,m_{F}=1)$. More details on the experimental setup are given in the  Appendix B. Information about initialization, quantum gates, and readout procedures are also given there.
@@ Acknowledgements
We acknowledge A. Ustinov for fruitful discussions and comments on the manuscript. The work was supported by Rosatom in the framework of the Russian Roadmap for Quantum computing (Contract No. 868-1.3-15/15-2021 dated October 5, 2021 and Contract No. 151/21-503 dated December 21, 2021). The work of N.A.N, A.S.N., and A.K.F. was supported by the Federal Academic Leadership Program Priority 2030, Strategic Project Quantum Internet (Project No. K1-2022-027 at MISIS). N.A.N. thanks the support of the Russian Science Foundation Grant No. 23-71-01095 (theoretical modeling of quantum circuits). E.O.K. thanks for the support of the Russian Science Foundation Grant No. 19-71-10091 (study of embedding qubits’ non-Hermitian dynamics into qutrits). The transmon-based device was supported by Rosatom and fabricated using the equipment of MIPT Shared Facilities Center (Contract No. 868/221-D dated October 24, 2022).
@@ A.1 General
fulfills the last equation in Eq. (18).
Thus, the key question is whether the first equation in Eq. (18) has a solution. In fact it does, provided
@@ A.3 Gate-level implementation
The last form emphasizes that the top left block of $R_{X}^{(12)}(\theta)$ reproduces $\Sigma/\sigma_{\rm max}$, while the unspecified entries * do not affect the resulting block encoding. They can be chosen arbitrarily (subject to the unitarity constraint), and $R_{X}^{(12)}(\theta)$ provides perhaps the simplest such choice.
@@ B.1 Initialization
Figure 6: Level scheme of the 171Yb+ ion. Solid lines show laser fields. Dashed lines show laser fields obtained by lasers phase modulation.
@@ B.3 Readout
The readout error sources include optical pumping from the $|0\rangle$ state to the ${}^{2}D_{3/2}(F=2)$ manifold during the transient stage in the beginning of each readout cycle, non-resonant pumping between qudit states, single-qudit gates infidelities and spontaneous decay of the ${}^{2}D_{3/2}(F=2)$ states Semenin et al. (2021). All these error contributions can be rather straightforwardly reduced by further optimization of the system parameters. For instance, it was demonstrated, that a duration of a single readout cycle in multiparticle processors based on 171Yb+ ions can be significantly decreased in comparison to our current results by reducing the amount of the stray light on the detector and increasing a photon collection efficiency Debnath (2016). This would significantly reduce errors due to spontaneous decay and non-resonant pumping.
@@ C.2.3 Readout
where $\langle$.$\rangle$ . ⟩ stands for the averaging over all trajectories and denotes the complex conjugation operation.